### v2

# Everybody Prune Now: Structured Pruning of LLMs with only Forward Passes

Abstract
Given the generational gap in available hardware between lay practitioners and the most endowed institutions, LLMs are becoming increasingly inaccessible as they grow in size. Whilst many approaches have been proposed to compress LLMs to make their resource consumption manageable, these methods themselves tend to be resource intensive, putting them out of the reach of the very user groups they target. In this work, we explore the problem of structured pruning of LLMs using only forward passes. We seek to empower practitioners to prune models so large that their available hardware has just enough memory to run inference. We develop Bonsai, a gradient-free, perturbative pruning method capable of delivering small, fast, and accurate pruned models. We observe that Bonsai outputs pruned models that (i) outperform those generated by more expensive gradient-based structured pruning methods, and (ii) are twice as fast (with comparable accuracy) as those generated by semi-structured pruning methods requiring comparable resources as Bonsai. We also leverage Bonsai to produce a new sub-2B model using a single A6000 that yields state-of-the-art performance on 4/6 tasks on the Huggingface Open LLM leaderboard.
Code for Bonsai can be found here
Machine Learning, ICML
[…]
1 Introduction
As large language models (LLMs) (OpenAI et al., 2023; Touvron et al., 2023; Gemini Team et al., 2023) continue to grow in size, the gap between models that achieve state-of-the-art performance and those that every-day machine learning (ML) practitioners can feasibly run on their available hardware continues to widen (Bender et al., 2021; Samsi et al., 2023). With the goal of democratizing access to these powerful models, previous research has proposed approaches such as pruning (Xia et al., 2022; Sun et al., 2023), distillation (Hinton et al., 2015; Gu et al., 2023) and quantization (Xiao et al., 2023) to create smaller models from larger pre-trained language models.
Figure 1: Perplexity versus inference speed-up of pruned models for methods that use only forward passes on the parent model. At any given target perplexity, Bonsai produces the fastest model, resulting in improved latency and throughput. Post-pruning adaptation is only possible for models below a certain size, given the available hardware. Circle sizes are proportional to the model’s memory footprint.
Table 1: Landscape of resource consumption (memory and compute) of different model compression methods at training time and the inference time resource consumption of the models they deliver. ✗  means the method incurs a prohibitive cost to the lay practitioner whilst ✓  denotes that it is a viable option with respect to that resource.
Regime Resource Approaches
Quantization (Mixed Precision)
Distillation
Unstructured Pruning
Gradient-Based Structured Pruning
Bonsai (Ours)
Train Memory
Compute
Inference Memory
Compute
Unfortunately, so far, these methods have fallen short of the goal of truly democratizing access to LLMs. In the process of producing a smaller model out of an LLM, methods such as distillation and structured pruning are inaccessible to the everyday practitioner due to their prohibitive resource consumption at training time. Specifically, pure distillation-based techniques require running LLMs to generate large amounts of teacher data (Jiao et al., 2019; Hsieh et al., 2023) whilst existing structured pruning approaches like LLM-Pruner (Ma et al., 2023) and LoRAPrune (Zhang et al., 2023) require several times more memory than is needed to run inference on the model being pruned. Though unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and quantization are less restrictive at training time, the models they produce are not faster except in the presence of specialized hardware for the former (Mishra et al., 2021), whilst the latter can actually slow down inference due to added overhead (Dettmers et al., 2022). This limits the usefulness of these options for practitioners who are concerned with latency-critical applications. Table 1 summarizes the landscape of existing methods and their limitations.
[…]
The practitioner only has enough memory on their hardware to run inference on the model to be pruned. [^1] Specifically we assume a forward pass with a batch size of at least 1. Given the hardware and open source LLM landscape as of January 2024, for an Nvidia A6000 (48GB) GPU, the largest (full-precision) model this would be applicable to is LLaMA-7B
The above setting is evergreen. As state-of-the-art models become more compute intensive over time, the generational gap in hardware available to the lay practitioner versus the most resource endowed institutions is expected to persist or possibly widen.
In light of the proposed setting, we present Bonsai, a forward pass-only structured pruning approach that is capable of delivering fast, compact, and accurate pruned models under the memory limitations that are typical of consumer hardware. To decide which modules (attention head, rows in feedforward projection, etc.) of the LLM to prune, Bonsai estimates module importances perturbatively by generating sub-models and evaluating their performance (running inference). We make this approach tractable by contributing multiple techniques.
First, we treat the problem of inferring each module’s importance from the performance of generated sub-models as an under-determined regression problem. This enables us to estimate the importance of a large number of modules by exploring a manageable number of random sub-models. This is unlike past perturbative approaches, which prohibitively require roughly as many sub-models as there are modules to select from (Ancona et al., 2020), making them intractable for LLMs. Next, instead of instantiating sub-models by dropping modules with equal likelihood (Kang et al., 2023), we use informative priors derived from work on unstructured pruning (Han et al., 2015; Sun et al., 2023). We thus obtain better estimates of module relevance with fewer evaluated sub-models. Finally, unlike past gradient-free approaches that greedily make pruning decisions layer-by-layer (Dekhovich et al., 2021; Nova et al., 2023; Sun et al., 2023), Bonsai takes a holistic view to preserve the accuracy of the pruned model: modules across layers are removed and evaluated together and relevance scores are computed globally to make pruning decisions. To the best of our knowledge, these ingredients taken together make Bonsai the first successful attempt at scaling gradient-free structured pruning to LLMs.
We conduct several experiments that demonstrate Bonsai’s efficacy. Bonsai, which uses only forward passes for pruning, achieves comparable performance to 2:4 semi-structured sparsity with Wanda (Sun et al., 2023) (0.75×0.75$\times$0.75 × speedup after fine-tuning with LoRA(Hu et al., 2021)) whilst also speeding up inference by up to 1.58×1.58$\times$1.58 × compared to the parent model. Even when compared to SoTA gradient-based structured pruning methods like LLM-Pruner and LoRAPrune, Bonsai outperforms these methods—on 4 out of 6 evaluation settings in our experiments. We also use Bonsai to prune the $\sim$3B Phi-2 (Li et al., 2023) model to a $\sim$1.8B model that outperforms the best sub-2B parameter model on the Huggingface Open LLM leaderboard on 4 out of 6 tasks as of the time of submission. Based on these strong results and its usability under real-world memory constraints, we view Bonsai as a significant contribution to unlocking the power of LLMs for a broader spectrum of practitioners facing diverse hardware constraints.
2 Background on Pruning, Problem Definition and Notation Setup
We are given an LLM, $\mathbf{M}_{\theta}$, parameterized by $\theta\in\mathbb{R}^{D}$. Also provided is $U$, a utility function that evaluates a model’s performance on a target task. We are interested in pruning $\mathbf{M}_{\theta}$ to produce a smaller and faster but still performant (with respect to $U$) model under the constraint that we only have enough memory to run inference on $\mathbf{M}_{\theta}$. Even though we assume we can run $\mathbf{M}_{\theta}$ on available hardware, pruning can be critical for achieving latency targets, reducing compute burden, or making the model small enough to adapt to new (out-of-domain) tasks by gradient-based fine-tuning.
Unstructured pruning approaches compress $\mathbf{M}_{\theta}$ by removing individual parameters $\theta_{j}$ from the model. This results in the updated model consisting of sparsified weight matrices with a smaller memory footprint. Unfortunately, the updated model does not enjoy inference speedups except when specialized hardware is available and thus poses a compute burden during inference. Whilst semi-structured variants – those that remove parameters in patterns like 2:4 or 4:8 (Mishra et al., 2021) – achieve some speedup, these are modest compared to those achieved with structured pruning.
Structured pruning takes a more modular view of the units to be removed from $\mathbf{M}_{\theta}$. Take that $\mathbf{M}_{\theta}$ is made up of modules $\mathbf{m}=\{m_{i}\}_{i\in[N]}$ each with corresponding parameter counts $\mathbf{s}=\{s_{i}\}_{i\in[N]}$ such that $\sum_{i}s_{i}=D$. For a transformer model, $\mathbf{m}$ could consist of attention heads, dimensions of fully connected layers, or even whole layers. For simplicity, we assume that $\mathbf{m}$ is made up of non-overlapping modules. Structured pruning compresses $\mathbf{M}_{\theta}$ by finding accurate sub-models defined by subsets of $\mathbf{m}$: provided with $\bar{\mathbf{m}}\subseteq\mathbf{m}$, we can construct an updated model $\mathbf{M}_{|\bar{\mathbf{m}}}$ that is produced by dropping the modules not in $\bar{\mathbf{m}}$ from $\mathbf{M}$. Thus, given a sparsity target $p$, structured pruning can be cast as the following combinatorial optimization problem:
$\begin{split}\mathbf{m}^{\ast}&=\mathrm{argmax}_{\bar{\mathbf{m}}\in\mathcal{F}_{p}}\quad U\left(\mathbf{M}_{|\bar{\mathbf{m}}}\right)\qquad\text{where}\\ \mathcal{F}_{p}&=\bigg{\{}\bar{\mathbf{m}}\subseteq\mathbf{m}~{}\bigg{|}~{}\bigg{(}\sum_{[j:m_{j}\in\bar{\mathbf{m}}]}s_{j}\bigg{)}\leq(1-p)D\bigg{\}}\end{split}$ (1)
$\mathcal{F}_{p}$ consists of all sub-models that meet the sparsity threshold. Note that, in general, not only does $\mathbf{M}_{|\mathbf{m}^{\ast}}$ have a smaller memory footprint than $\mathbf{M}$, it is also faster to run inference on it since it has fewer modules.
Many structured pruning methods attempt to solve Equation 1 by gradient-guided optimization (or search) over the space of sub-models. However, since we are interested in the memory-constrained setting where computing gradients is not feasible, these methods cannot be used. We will thus focus on developing a memory-friendly structured pruning technique, but we will compare to semi-structured pruning methods since they also have minimal memory overhead.
3 Methodology
In this section, we will develop a structured pruning algorithm that relies exclusively on forward passes through the parent model.
3.1 Estimating module relevance with only forward passes
We have motivated the setting of pruning a model that is so large (relative to the amount of memory available to the practitioner), such that we can only run forward passes through it. This means that we have to solve Equation 1 by relying on only evaluations of $U$, as opposed to gradient-based optimization with respect to $U$. A brute force approach would enumerate all members of $\mathcal{F}_{p}$, find their corresponding performances and pick the best one. Unfortunately, this would be computationally infeasible since $|\mathcal{F}_{p}|$ is combinatorial in the size of the model. [^2] The typical fully connected (FC) sub-layer in an LLM has size $>10^{4}$ units. Pruning just 1 of these FCs to 0.5 sparsity in this way requires evaluating over $\approx$ ($10^{4}\choose 10^{3}$ subsets.
We propose a computationally tractable approach where we first perform a small number, $n$, of evaluations, where $n\ll|\mathcal{F}_{p}|$, to gather data for estimating the relevance of each module in $\mathbf{M}$ with respect to the metric $U$. Upon carrying out this estimation, we can greedily choose the member of $\mathcal{F}_{p}$ that has the highest total module relevance. Specifically, let us assume that we have estimated $\mathbf{\beta}=\{\beta_{i}\}_{i\in[N]}$ to be the relevance of each of the $N$ modules. We can generate an approximate solution to Equation 1 as:
$\mathbf{m}^{\ast}\approx\mathbf{m}^{\mathrm{approx}}=\mathrm{argmax}_{\bar{\mathbf{m}}\in\mathcal{F}_{p}}\quad\sum_{j\in\bar{\mathbf{m}}}\beta_{j}$ (2)
We note that Equation 2 is straightforward to solve, as it simply requires sorting $\beta_{j}$s and greedily selecting the top modules until the parameter count constraint is met (this greedy approach may slightly overshoot the sparsity constraint but since $s_{i}\ll(1-p)D~{}\forall~{}i$ for our settings of interest, the difference is not significant).
Estimating $\mathbf{\beta}$:
To obtain estimates of the module relevance scores $\mathbf{\beta}=\{\beta_{i}\}_{i\in[N]}$, we propose to generate and evaluate $n\ll|\mathcal{F}_{p}|$ sub-models $\bar{\mathbf{m}}_{k}$, and construct a dataset of the sampled sub-models and their corresponding evaluated performances: $\mathbb{D}=\left\{\bar{\mathbf{m}}_{k},U_{k}\right\}_{k\in[n]}$ where $U_{k}=U(\mathbf{M}_{|\bar{\mathbf{m}}_{k}})$. Given $\mathbb{D}$, we treat the problem of estimating $\mathbf{\beta}$ as an under-specified regression problem :
$\begin{split}\hat{\mathbf{\beta}}&=\mathrm{argmin}_{\mathbf{\beta}\in\mathbb{R}^{N}}\quad\frac{1}{n}\sum_{\left(\bar{\mathbf{m}}_{k},U_{k}\right)\in\mathbb{D}}\left(U_{k}-\beta^{T}\alpha_{\bar{\mathbf{m}}_{k}}\right)^{2}+\gamma\|\mathbf{\beta}\|\\ \end{split}$ (3)
where $\alpha_{\bar{\mathbf{m}}_{k}}\in\mathbb{R}^{N}~{}\big{|}~{}\left(\alpha_{\bar{\mathbf{m}}_{k}}\right)_{i}=\mathbf{1}[i\in\bar{\mathbf{m}}_{k}]$, is the binary vector that has 0 at indices where modules have been dropped.
Implementing a sub-model $\bar{\mathbf{m}}_{k}$ as a binary mask $\alpha_{\mathbf{m}_{k}}$ is key to practically realizing our approach. We never actually fully instantiate sub-models as this would be prohibitively expensive. Instead, we create sub-models virtually by zeroing out the outputs of the components to be pruned so they have no effect on the final model output.
3.2 Selecting sub-models for evaluation
An as yet unexplored design choice is how to choose the $n$ candidate sub-models for evaluation. A naive approach would be to sample uniformly from the space of all feasible sub-models. However, we can show that this is sub-optimal. Let us assume $m_{i}$ is a module that is critical for good performance under evaluation with $U$. Since $n<N$, it means that some modules will never be ”turned on” in the list of $n$ chosen sub-models. If $m_{i}$ happens to be one of these masked-out modules, the resulting estimate of $\hat{\beta_{i}}=0$ would in turn result in $\mathbf{m}^{\mathrm{approx}}$ being a poor estimate for $\mathbf{m}^{\ast}$. We therefore need to make a more informed choice to ensure that our estimates of $\mathbf{\beta}$ are accurate and useful.
Given a module $m_{i}$, we can set the likelihood of it being present in any of the $n$ sampled sub-models to be proportional to some measure $\rho_{i}$ which captures our prior belief that $m_{i}$ is a useful module. To define $\rho_{i}$, we can turn to metrics from the unstructured pruning literature. Specifically, we set $\rho_{i}$ to be a module-level analogue of any of the parameter-level unstructured pruning metrics like Wanda (Sun et al., 2023) or output activation magnitude. As an illustrative example, consider that we have shallow network with 1 hidden layer (of dimension $d$) we wish to prune. If output activation magnitude is the metric for our prior, we would calculate the activation vector for the $d$ hidden dimensions – $\hat{\mathbf{a}}$ – averaged across multiple samples. The likelihood that the $i$th column of the matrix $W\in R^{d\times d}$ is left unpruned would be:
$\rho_{i}\propto\hat{\mathbf{a}}_{i}=\frac{1}{B}\sum_{b}\bigg{|}\sigma\bigg{(}\big{(}W^{T}[i,:]\big{)}x_{b}\bigg{)}\bigg{|}$
$\sigma\text{ is the nonlinearity}$. Sampling sub-models according to $\mathbf{\rho}$ means we are more likely to explore models that already have a good performance $U$. The priors $\rho_{i}$ can be computed via forward passes through the original unmodified model $\mathbf{M}_{\theta}$ which allows us to respect memory constraints.
To enhance the efficiency of our method, instead of considering all active modules for pruning, in each layer we consider the bottom $2p$ fraction of modules [^3] The choice of $2p$ is arbitrary. Practitioners can run hyper-parameter searches to find more optimal values for their settings. ranked according to our prior $\rho$. Thus, the top $1-2p$ fraction of modules remain fixed whilst the rest are perturbed and their relevance scores are computed and compared for pruning. This helps reduce the space of possible sub-models being considered for evaluation. For the bottom $2p$ fraction being compared, whenever we generate a mask $\alpha_{\mathbf{m}_{k}}$ with sparsity level $p$, we also generate its complement $\alpha^{c}_{\mathbf{m}_{k}}$, which is obtained by flipping the values in $\alpha_{\mathbf{m}_{k}}$ (except for the fixed $1-2p$ fraction of entries). Covert & Lee (2020) show that this technique can help reduce the variance of the estimator obtained from regression with binary inputs.
Algorithm 1 Bonsai
1:  Input: Model [$\mathbf{M}_{\theta}$], sub-models per iteration [$n_{\mathrm{iter}}$] Sparsity per iteration [$p_{\mathrm{iter}}$], Target sparsity [$p$] Module list [$\mathbf{m}$]
2:  for $l=1$ to $\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ do
3:     $\mathbb{\rho}^{l}\leftarrow$ Calculate unstructured pruning metric for all modules in $\mathbf{m}$
4:     $\bar{\mathbb{\rho}}^{l}\leftarrow$ Fix the top $(1-2p_{\mathrm{iter}})$ of $\mathbb{\rho}^{l}$ to $\infty$
5:     Sample $\{\mathbf{\bar{m}}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$ sub-models according to $\bar{\mathbb{\rho}}^{l}$
6:     Generate $\{\mathbf{\bar{m}}^{c}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$ complements of $\{\mathbf{\bar{m}}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$
7:     Run forward pass on each sub-model and compute $U$. Construct $\mathbb{D}^{l}=\left\{\bar{\mathbf{m}}_{i},U_{i}\right\}_{[\frac{n_{\mathrm{iter}}}{2}]}\bigcup\left\{\bar{\mathbf{m}}^{c}_{k},U_{k}\right\}_{[\frac{n_{\mathrm{iter}}}{2}]}$
8:     $\beta^{\l}\leftarrow\mathrm{Regress}\left(\mathbb{D}^{l}\right)$
9:     $\{m_{\mathrm{pruned}}\}\leftarrow$ sort $\beta^{\l}$ and remove the bottom $k$ modules that correspond to $p_{\mathrm{iter}}$ fraction of the model.
10:     $\mathbf{m}\leftarrow$ update module list to exclude $\{m_{\mathrm{pruned}}\}$
11:  end for
12:  Output: Pruned model $\mathbf{M}_{|\mathbf{m}}$
3.3 Iterated Pruning
Previous work on gradient-based pruning (Anwar et al., 2017; Frankle & Carbin, 2018) have shown that taking an iterated approach to pruning yields improved results over pruning directly to the target sparsity $p$. Similarly, we define an updated pruning fraction $p_{\mathrm{iter}}<p$ and perform $\mathrm{iter}=\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ steps where we explore $n_{\mathrm{iter}}=\lceil\frac{n}{\mathrm{iter}}\rceil$ sub-models at a time. At the beginning of each iteration, we re-estimate the priors $\mathbf{\rho}$ for the unpruned modules and use them to select the $n_{\mathrm{iter}}$ sub-models to evaluate.
We combine the recipes developed in Sections 3.1, 3.2 and 3.3 together to produce Bonsai [^4] Structural pruning is a canonical way of giving a bonsai tree its shape hence the name., our gradient-free structural pruning algorithm. Algorithm 1 specifies Bonsai in detail.
4 Post-pruning adaptation
Depending on the sparsity level $p$, it is possible to obtain a pruned model on which it is feasible to run a parameter-efficient finetuning method like LoRA (Hu et al., 2021) with the available hardware memory. In this case, we can fine-tune the pruned model (result of Algorithm 1) on the downstream task in order to recover some of the parent model performance that was degraded by pruning.
Like many past works (Sanh et al., 2020; Xia et al., 2022), we combine pruning with distillation by incorporating a distillation loss in the training objective during fine-tuning of the pruned model. Let $\mathcal{L}_{\mathrm{task}}$ be the loss function over the task data and $\mathcal{L}_{\mathrm{distill}}$ be the distillation objective. We optimize the following post-pruning objective: $\mathcal{L}^{\mathrm{post-prune}}=\mathcal{L}_{\mathrm{task}}+\lambda\mathcal{L}_{\mathrm{distill}}$. Using $i$ to index the task data, we have:
$\mathcal{L}_{\mathrm{distill}}=\sum_{i}D_{\mathrm{KL}}\bigg{(}\mathrm{logits}^{i}\left(\mathbf{M}_{|\mathbf{m}^{\mathrm{approx}}}\right)~{}\|~{}\mathrm{logits}^{i}\left({\mathbf{M}}\right)\bigg{)}$, where $\lambda$ is a scalar weighting that can be optimized via cross-validation. Also, distillation can be performed without significant memory overhead by apriori caching the logits from the parent model $\mathbf{M}$ instead of hosting the model in memory during fine-tuning.
Table 2: Wikitext-2 perplexity for 50% sparsity of LLaMA-2 7B with end-to-end latency speedups (relative to LLaMA-2 7B). We use Phi-2 (Li et al., 2023) at the target size as a strong model for comparison. For semi-structured (Wanda 2:4) pruning, after fine-tuning, the learned LoRA weights cannot be merged with the primary model weights else the model reverts back to being dense; this leads to the reported slow downs. All methods use forward passes only on the parent model.
[…]
Phi-2 3B $8.69$ 1.24×1.24$\times$1.24 ×
LLaMA-2 7B Pruned
Wanda 2:4 3B $10.52$ 1.14×1.14$\times$1.14 ×
$8.34$ 0.75$\times$
Bonsai 3B $8.89$ $\boldsymbol{1.58}\times$
5 Experimental Details
All experiments are conducted on a single NVIDIA A6000 (48Gb) GPU. In all experiments with Bonsai, we prune (1) the heads in the self-attention layers (2) the dimensions of the fully connected layers.
Comparing Forward Pass Only Methods: We focus our first set of experiments on comparing methods that can be run without gradient-based optimization. We consider pruning the LLaMA-2 7B model (Touvron et al., 2023) to 50% sparsity. We evaluate on the Wikitext-2 (Merity et al., 2016) validation dataset and so our signal for pruning, $U$, is the language modelling performance on the training set. When measuring speedups, we consider end-to-end latency of running inference on model.sequence_length chunks of the Wikitext-2 validation set. See Appendix A.2 for details about the hyper-parameters used for all methods in this experiment and for specifics about fine-tuning.
Comparing Structured Pruning Approaches: We compare Bonsai to the following gradient-based structured pruning approaches: LLM-Pruner (Ma et al., 2023) and LoRA-Prune (Zhang et al., 2023). We prune the LLaMA-1 7B model (Touvron et al., 2023) to 50% sparsity since this was the model version available at time of release of the above methods. We compare these methods on Wikitext-2 and also on 6 tasks from the Eleuther LLM Evaluation Harness (Gao et al., 2023). Pruning signal for Wikitext-2 task is the same as the above experiment. For the Eleuther Harness tasks, we use language modelling performance on the C4 (Raffel et al., 2020) dataset as pruning signal. We also do parameter efficient finetuning on our pruned model with 30K 512-length sequences from this corpus. Specific hyper-parameters for Bonsai for this experiment can be found in Appendix A.3.
6 Main Results
6.1 Bonsai produces fast and performant models with only forward passes
In Table 2, we explore options that are available to practitioners when they can only run forward passes on the parent model. Here, LLaMA-2 7B is compressed to 50% sparsity.
We compare Bonsai to the semi-structured variant of Wanda (Sun et al., 2023). Before fine-tuning, the model produced by Wanda 2:4 achieves a speedup over the parent model (1.14$\times$) but is slower and less accurate than Bonsai’s pruned model (8.89ppl vs 10.52ppl). While the performance gap can be bridged by LoRA finetuning (10.52 $\rightarrow$ 8.34), the adapted semi-structured model experiences a drastic slowdown (0.75$\times$), since the learned low-rank matrices cannot be merged with the original sparsified ones without reverting back to dense computation. Thus, LoRA fine-tuned Wanda 2:4 is twice slower (0.48$\sim$ 0.48$\times$ 0.48 ×) than the model from Bonsai.
In this memory-constrained setting, practitioners could alternatively opt for a pre-existing model of the target size instead of pruning a larger model. We compare the Bonsai-pruned model to Phi-2 (Li et al., 2023), a strong representative pre-existing model of similar size. As can be seen in Table 2, Bonsai is able to generate a model that is as accurate (0.2 difference in ppl) yet significantly faster (1.58$\times$ vs. 1.24$\times$ speedup), thus making it a competitive option to consider even if a model already exists at the target size.
Table 3: LLaMA-1 (50% sparsity) after post-pruning adaptation with LoRA. ${}^{\dagger}$ indicate results as reported by Zhang et al. (2023).
Method Foward-only Wikitext-2 $\downarrow$ BoolQ HellaSwag WinoGrande ARC-e ARC-c Average $\uparrow$
LLaMA1-7B (Touvron et al., 2023) - 5.68 75.05 56.92 69.93 75.34 41.89 63.83
${}^{\dagger}$ (Ma et al., 2023) 16.41 60.28 47.06 53.43 45.96 29.18 47.18
${}^{\dagger}$ (Zhang et al., 2023) 11.60 61.88 47.86 55.01 45.13 31.62 48.30
Bonsai 10.92 67.22 43.09 61.64 54.92 26.28 50.63
6.2 Bonsai outperforms other structured pruning approaches
We next investigate the setting where the practitioner has enough memory to run gradient-based structured pruning methods. We compare Bonsai to recent SoTA methods: LLM-Pruner (Ma et al., 2023) and LoRA-Prune (Zhang et al., 2023). Since these approaches report their results for the LLaMA-1 only, we prune LLaMA-1 7B to 50% sparsity.
As evinced by Table 3, Bonsai outperforms existing structured pruning methods for LLMs even though it exclusively uses forward passes in the pruning stage. We attribute the superior performance of Bonsai to the fact that its pruning decisions are informed by directly exploring the space of sub-models whilst the other approaches resort on inaccurate proxies of module relevance in order to reduce the memory overhead of a fully gradient-based optimization approach. Specifically, LoRA-Prune’s criterion for deciding which modules to prune uses gradient signals not from the modules directly but from a surrogate low-rank matrix. On the other hand, LLM-Pruner attempts to model the impact of removing a module on the model loss via a Taylor expansion and substitutes the LLM’s intractable (memory-wise) Hessian term with the Fisher information matrix which we posit leads to sub-optimal pruning decisions.
6.3 Bonsai can produce compressed models with strong zero-shot abilities
Table 4: Phi-2 pruned to 35% sparsity compared to the best sub-2B parameter models on the Hugging Face OpenLLM Leaderboard. At the time of submission, $\dagger$ was the best sub-2B parameter model on the Leaderboard.
Generation Multiple Choice (MC)
Model Size GSM8k (5-shot) ARC-c (25-shot) Winogrande (5-shot) Hellaswag (10-shot) Truthful-QA (0-shot) MMLU (5-shot) MC Average $\uparrow$
Phi-2 (Li et al., 2023) 2.7B 54.81 61.09 74.35 75.11 44.47 58.11 62.63
StableLM-2-1_6$\dagger$ 1.6B $\boldsymbol{17.44}$ 43.34 64.56 $\boldsymbol{70.45}$ 36.78 38.95 50.82
Ours 1.8B 6.37 $\boldsymbol{47.44}$ $\boldsymbol{68.35}$ 65.09 $\boldsymbol{42.20}$ $\boldsymbol{40.53}$ $\boldsymbol{52.72}$
Considerable amounts of compute and data, beyond what is available to lay practitioners, are needed to train LLMs with strong zero-shot capabilities (OpenAI et al., 2023; Gemini Team et al., 2023). In this section, we demonstrate that Bonsai can empower everyday practitioners to produce strong and compact models with competitive zero-shot abilities by simply pruning bigger models on their available hardware.
We use Bonsai to prune a $\approx$3B Phi-2 model to $\approx$1.8B (35% sparsity). Values for the Bonsai hyper-parameters in this experiment are in Appendix A.4. Since its relatively small, the 1.8B pruned model can be fully fine-tuned on 1 A6000 GPU over 100k sequences of 2,048 tokens from the C4 dataset. As can be seen from Table 4, our pruned model achieves strong zero-shot performance compared to StableLM, the leading sub-2B parameter LLM on the Hugging Face OpenLLM leaderboard (Gao et al., 2023) — outperforming it on 4 out of 6 tasks as of January 2024. Interestingly, one exception to the general trend of Bonsai’s superior performance is the GSM-8K dataset, which is a mathematical reasoning dataset that requires generation of a long reasoning chains. We posit that this is because currently, Bonsai prunes with respect to language modeling likelihood, as opposed to reasoning accuracy. An interesting avenue for future work is to prune to improve maintain reasoning ability.
7 Analysis
In this section, we conduct various ablative experiments to understand how the methodological ingredients from Section 3 contribute to make Bonsai effective.
Do we need both perturbative and regressive components of Bonsai? Figure 2 shows that both components are key to obtaining a good pruned model. Removing the estimation of module importances via regression leads to a degradation in performance ($61.6$ ppl $\rightarrow$ $146.6$ ppl). More degradation ($146.6$ ppl $\rightarrow$ $405.7$ ppl) is encountered if we do not perform any perturbative evaluations on the parent model but simply just prune according to the prior $\rho$ as computed from the unperturbed parent model.
Figure 2: LLaMA-2 7B pruned to 50% sparsity. Perturbatively evaluating the impact of removing modules and then using regression to estimate module importances are both key in making Bonsai effective. Details of experiment configuration can be found in Appendix B. No post-pruning adaptation is performed.
Table 5: Varying the number of perturbative evaluations. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix G for details. No post-pruning adaptation is performed.
$\mathrm{ns}=50$ $\mathrm{ns}=200$ $\mathrm{ns}=1000$
PPL ($\downarrow$) NaN $61.63$ $22.09$
How many perturbative samples are sufficient? We investigate the number of perturbative samples required to obtain good estimates of module importances after performing regression as in Equation 3. Our results are shown in Table 5. As expected, performance improves as we increase the number of sub-models explored. We note that the number of samples we explore, $\mathrm{ns}$, is significantly less than the number of candidate modules at each iteration ($N\approx 70k$), nevertheless Bonsai is able to deliver a performant pruned model because of the recipes developed in Section 3.
How much performance is recovered by post-pruning adaptation? During iterative pruning, Bonsai damages the parent model by removing modules but does not perform intermittent retraining to recover lost performance since even intermediate models may be too large for fine-tuning. Even so, as Table 6 shows, the final model produced by Bonsai has reasonable performance without fine-tuning. We attribute this to the general robustness of LLMs, the redundancy of modules with respect to target end-tasks and Bonsai’s ability to identify good candidates for pruning. If the pruned model is small enough in size, we can perform either full fine-tuning or parameter-efficient fine-tunin to recover more performance, as can be seen from Table 6.
Table 6: Impact of post-pruning adaptation of LLaMA-2 7B pruned to 50% sparsity. See Appendix E for more details.
Method Wikitext-2 PPL
No Post-Pruning Adaptation $19.47$
Post-Pruning Finetuning $10.39$
   +  Distillation $8.89$
Figure 3: LLaMA-2 7B pruned to 50% sparsity. See Appendix F for details of experiment configuration and definitions of module-level analogues of Wanda and Activation Magnitude.
What is the impact of the choice of metric for the prior $\rho$? We investigate three different choices of metrics for defining $\rho$. Figure 3 shows that using the module-level analogue of Wanda (Sun et al., 2023) yields the best performance, both before and after post-pruning adaptation. This indicates that Wanda is a strong signal for efficiently estimating the importance of model units (whether at the parameter- or module-level).
Should Bonsai prune iteratively? Table 7 demonstrates the benefits of using Bonsai  in an iterative fashion. Pruning slowly ($p_{\mathrm{iter}}=0.05$) yields the best results, but this comes at the cost of increasing the total time to prune the model. The performance gap between values of $p_{\mathrm{iter}}$ persists even after post-pruning adaptation, indicating that slower pruning allows for more accurate estimates of module importance.
Table 7: Varying $p_{\mathrm{iter}}$. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix C for experiment details.
$p_{\mathrm{iter}}=0.05$ $p_{\mathrm{iter}}=0.1$ $p_{\mathrm{iter}}=0.2$
w/o Adapt $19.47$ $61.63$ $209.44$
w Adapt $8.89$ $9.15$ $9.57$
8 Related Work
8.1 Pruning
Unstructured pruning:
Whilst structured pruning approaches remove entire model components like layers (Xu et al., 2020; Xia et al., 2022), dimensions of linear layers (Wang et al., 2019) or attention heads (Michel et al., 2019; Held & Yang, 2022), unstructured pruning (Han et al., 2015; Frankle & Carbin, 2018; Benbaki et al., 2023; Sun et al., 2023) removes individual parameters of the model. These approaches achieve memory savings by inducing sparsity in the model weights, but they generally do not result in actual model speedups except when specialized hardware is available (Mishra et al., 2021). Proposed semi-structured sparsity methods (Mishra et al., 2021) such as 2:4 and 4:8 patterns do result in faster inference, but the speedup gains they achieve are far from the idealized 2×2$\times$2 ×. There are several gradient-free, unstructured pruning approaches. Since, as far as we know, there are no scalable gradient-free structured pruning alternatives outside of Bonsai, we compare it with Wanda (Sun et al., 2023), a SoTA unstructured/semi-structured pruning technique for LLMs.
Structured pruning without backward passes:
To the best of our knowledge, all existing structured pruning techniques for large (over 1B scale) language models, like LoRAPrune (Zhang et al., 2023) and LLM-Pruner (Ma et al., 2023), are gradient-based. Thus, under the memory setting we consider, these methods cannot be applied since their memory requirements well exceed the memory needed for inference on the model to be pruned. For smaller language models like BERT, Nova et al. (2023) recently proposed Kernelized Convex Masking (KCM) for gradient-free structured pruning. Unfortunately, to prune a fully connected layer with $K$ intermediate units, KCM requires instantiating a $K\times K$ coefficient matrix (for each layer). Whilst this is feasible for BERT models, a typical LLM (OpenAI et al., 2023; Touvron et al., 2023) has $K\sim 10^{4}-10^{5}$ which would make the size of each per-layer coefficient matrix comparable to the size of the LLM itself. In computer vision, there exist several perturbative gradient-free structured pruning techniques (Ancona et al., 2020; Dekhovich et al., 2021). However, these methods have been exclusively applied to small vision models (VGG; (Simonyan & Zisserman, 2014), WideResnet; (Zagoruyko & Komodakis, 2016)) and, taken as is, would not scale to LLMs. We introduce several techniques to make Bonsai scale to large language models. A key difference between Bonsai and these perturbative methods is that pruning decisions are made globally instead of on a local layer-by-layer basis (Ancona et al., 2020; Dekhovich et al., 2021). A failure mode of layer-by-layer pruning is that if a module, say $m_{i}$, which is in the top $p$ fraction of relevant modules when the whole model is considered but in the bottom $1-p$ fraction within its layer will not be included in the pruned model — a clearly sub-optimal choice. Bonsai enables us to obtain globally meaningful estimates of module relevance that are comparable across layers.
8.2 Model compression beyond pruning
Distillation:
Originally proposed by Hinton et al. (2015), Sanh et al. (2019) have shown that it is possible to reach similar performances on many downstream tasks using much smaller language models pre-trained with knowledge distillation. The resulting models are lighter and faster at inference time. Unfortunately, as noted by Xia et al. (2022), distilling a pre-trained model to a student model that has been initialized from scratch requires lots of compute and data to recover reasonable levels of teacher performance. To reduce this resource burden, Sanh et al. (2020); Lagunas et al. (2021); Xia et al. (2022) combine distillation with gradient-based structured pruning, so that the student model does not have to be initialized from scratch. Bonsai also leverages distillation, but only during the post-pruning adaptation phase. This maintains the constraint that the pruning process is gradient-free. And since the pruned model is smaller than the parent model, we can safely perform distillation and parameter efficient fine-tuning (Hu et al., 2021; He et al., 2021) within the bounds of available memory.
Quantization:
Li et al. (2020) introduced a training strategy that showed that heavily compressed, large models achieved higher accuracy than lightly compressed, small models, providing both efficiency and high-accuracy results on NLP tasks at once. Dettmers et al. (2022) introduced a new method called ‘LLM.int8()‘, which allows loading large models with 16 or 32-bit weights and using them immediately for inference without any performance degradation by combining vector-wise quantization with mixed-precision decomposition. Xiao et al. (2023) enabled 8-bit weight, 8-bit activation (W8A8) quantization for LLMs with >100B parameters, which smoothed activation outliers by migrating the quantization difficulty from activations to weights with a mathematically equivalent transformation. Quantization tends to be complementary to pruning, allowing independent exploration and subsequent combinations of both techniques. Some quantization approaches, specifically mixed precision ones like Dettmers et al. (2022) (Appendix D), can produce resulting models that are slower than the parent model, though it has also been shown that it can result in more accurate models than pruning (Kuzmin et al., 2023). This paper focuses on gradient-free structured pruning. We leave combining Bonsai with quantization for future work.
9 Conclusion, Limitations, and Future Work
In this work, we have presented Bonsai, the first tractable gradient-free method for structured pruning of LLMs. Bonsai allows practitioners to prune any LLM as long as they have enough memory to run inferences on the model. Through a battery of experiments, we have shown that the models produced by Bonsai are small, fast, and accurate.
A primary limitation of Bonsai is its runtime. As Section 7 shows, performance improves when we increase the number of sub-models explored, slow down pruning, and use more data samples; but this comes at the cost of increased runtime. Our result in Table 2 took $\sim 40$ hours to obtain, whilst Wanda-2:4 takes on the order of minutes. Since the model we produce is twice as fast, this trade-off is worthwhile when amortized over inference on the pruned models.
Bonsai presents several avenues for expansion by future work. First, though sub-models are sampled from an informative prior, $\rho$, the sampling process is not adaptive. Bonsai could further be strengthened by dynamically exploring the space of sub-models. Next, because of memory constraints, Bonsai does not fine-tune the model during iterative pruning to recover degraded performance. However, forward-pass-only fine-tuning approaches like MeZO (Malladi et al., 2023) exist that can be used to dynamically update the model during pruning.
10 Acknowlegements
This work was supported in part by the National Science Foundation grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, and funding from Meta, Morgan Stanley, Amazon, and Google. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies. We are grateful for helpful feedback from Mingjie Sun, Victor Akinwande, Asher Trockman, Afshin Rostamizadeh and Daniel Glasner
11 Broader Impact Statement
Network pruning is an effective compression strategy to reduce the inference costs of large language models. However, existing techniques for structured pruning impose expensive memory constraints that may render them out of reach in practice. By reducing such constraints, work aims to democratize pruning to provide practitioners the ability to produce their own pruned LLMs for real-world applications. While we ultimately view this democratization as a benefit, we note that a necessary outcome is that it may also make it more feasible for malicious actors to readily use LLMs in practice. Finally, we note that our work has primarily focused on the axes of efficiency and utility (e.g., perplexity) in assessing performance; recent work has shown that compression may also have outsized effects on issues such as model fairness and robustness, which would be interesting additional aspects to consider in future study.
Appendix A Main Experiment Details
A.1 Hyper-parameters for all Bonsai regression during pruning
When using Bonsai, we estimate $\beta$ from $\mathbb{D}^{l}$ by performing linear regression via gradient descent with Adam (Kingma & Ba, 2014). We cross-validate over the following set of hyper-parameters. Note that doing this cross-validation takes much less time than the time needed to construct the dataset $\mathbb{D}^{l}$.
Table 8: Bonsai hyper-parameters for regression. This applies to all experiments unless otherwise specified
$\gamma$(Regression Weight) Learning rate Batch Size Epochs
{100, 0, 1e-4} {100, 10, 1, 0.1} {32, 64, 128} 50
During cross validation, we choose the model whose predictions have the best Kendall rank correlation co-efficient (Kendall, 1948) with the target. We do this because we do not care about matching $U_{k}$ exactly for each sub-model $k$; we rather care that our learned $\beta$ predicts the correct rankings amounts sub-models, which would denote that $\beta$ reasonably models relative module importances.
In general, we use $\ell_{1}$-norm regularization on $\beta$ for all experiments. For the Phi-2 experiment in Section 6.3, we find that $\ell_{2}$-norm works better.
A.2 Forward Pass Only Experiments
Table 10 show the Bonsai hyperparameters we used for the experiments in Section 6.1.
Table 9: Bonsai hyper-params for forward only experiments
Table 10: Bonsai  fine-tuning HP for pruned LLaMA family models
$p_{\mathrm{iter}}$ $\mathrm{ns}_{\mathrm{sub-models}}$ $\mathrm{ns}_{\mathrm{data}}$ Metric for $\rho$
0.05 200 32 (per-iter) Wanda
LR rank LoRA-$\alpha$ $\lambda$ (Distill Weight) LoRA Modules
1e-4 128 4$\times$×rank 0.01 All Modules
Table 10: Bonsai  fine-tuning HP for pruned LLaMA family models
For Wanda(Sun et al., 2023), we use the default hyper-parameters specified by the paper repo here for pruning. For fine-tuning, we use rank = 64. We apply LoRA to only the q_proj and v_proj matrices in each layer of the pruned LLaMA model – this is unlike with Bonsai where we fine-tune all modules. We cannot do same because since the Wanda model just produces sparse matrices, the matrices instantiated during the backward pass are the same sizes as the sparsified matrices and thus occupy more memory (compared to our approach that actually makes the matrices smaller in dimension instead of sparsifying). We are also unable to perform distillation on the Wanda models due to this reason. For fine-tuning the Phi-2 model on Wikitext-2, we use the same hyper-parameters as Bonsai in Table 10.
A.3 Experiments comparing to Gradient based structured pruning
We compare to LoRA-Prune and LLM-Pruner. We take their performance results directly from the LoRA-Prune paper. Whilst we use 1 A6000 GPU (48G) for all experiments, LoRA-Prune uses A100 GPU (80G) for pruning LLaMA-1 7B.
All Bonsai hyper-parameters are the same as Appendix A.2 except for $\mathrm{ns}_{\mathrm{sub-models}}$ which we set to $1000$.
A.4 Phi-2 pruning experiment details
For the experiment in Section 6.3, All Bonsai hyper-parameters are the same as Appendix A.2 except for the following changes:
  • $\mathrm{ns}_{\mathrm{sub-models}}=2000$
  • $p_{\mathrm{iter}}=0.35$. We thus perform 1-shot pruning directly to the target sparsity of 35%. We find that this seems to work best for the Phi-2 model. We posit that this might be because the Phi-2 models use LayerNorm(Ba et al., 2016) whilst the other models we explore, LLaMA and Mistral use RMSNorm.
  • Due to its relatively small size, the 1.8B pruned model can be fully fine-tuned on a single A6000 GPU over 100k sequences of length 2,048 tokens from the C4 dataset instead of using LoRA.
Appendix B Impact of regression and perturbation ablation details
For the experiment in Section 3, All Bonsai hyper-parameters are the same as Appendix A.2 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
A simple alternative to Bonsai is to leverage the prior $\rho$, computed from the unperturbed parent model, and make pruning decisions exclusively according to this. This is the No Perturbation + No Regression baseline in Figure 2. This approach has quite poor performance. We can further improve this baseline by adding back perturbative aspect where we prune the parent model according to $\rho^{\prime}$ which is computed by aggregating the $\rho$ metric computed over the perturbed models. Note that we use a Wanda based metric to define $\rho$ for this experiment. Module level analogues of the unstructured pruning metrics we explore are defined in Appendix F.
Table 11: Experiment on linear regression to estimate module importances. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
Linear Regression Relative Speepdup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
No $2.06$ $146.57$ $9.68$
Yes $1.77$ $61.63$ $9.15$
Appendix C Varying the pruning fraction per-iteration
For the experiment in Section 3, All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $p_{\mathrm{iter}}$.
Table 12: Varying the fraction pruned at a time. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
Prune Frac Relative Speepdup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
0.05 $1.58$ $19.47$ $8.89$
0.1 $1.77$ $61.63$ $9.15$
0.20 $1.67$ $209.44$ $9.57$
Appendix D Varying the number of calibration data points for pruning
All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\mathrm{ns}_{\mathrm{data}}$ and $p_{\mathrm{iter}}=0.1$ to speed up pruning.
Table 13: How many data-points to consider during forward passes. Wikitext-2 Perplexity. Llama-2 7B pruned to 50% sparsity
$\mathrm{ns}_{\mathrm{data}}$ w/o Adapt w Adapt
8 $130.04$ $9.45$
32 $61.63$ $9.15$
Appendix E Post-pruning adaptation
For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2.
Appendix F Impact of prior
For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\rho$
F.1 $\rho$ is Activation Magnitude
MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned. Note that for all transformer models evaluate, the MLP components are 2 layer and thus have a single intermediate dimension. For any data-sample sequence $b$, we flatten model activation at this point $\mathbf{a}\in\mathbb{R}^{B\times S\times d}\rightarrow\mathbb{R}^{BS\times d}$ and then compute the following averaged activation magnitude :
$\left(\rho\in\mathbb{R}^{d}\right)\propto\hat{\mathbf{a}}=\frac{1}{B}\sum_{b}\mathrm{Mean}\bigg{(}\big{|}\mathbf{a}_{b}\big{|},\mathrm{axis=}0\bigg{)}$ (4)
Self-Attention Module: For any data-sample sequence $b$, the output of the self attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output. We can flatten $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}\rightarrow\mathbb{R}^{BSd_{h}\times h}$ and then use the same formula as Equation 5 above to calculate $\rho$.
$\left(\rho\in\mathbb{R}^{h}\right)\propto\hat{\mathbf{a}}$ (5)
F.2 $\rho$ is Wanda (Sun et al., 2023)
MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned. Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the MLP. For any data-sample sequence $b$, we flatten model activation before the final output, $\mathbf{a}\in\mathbb{R}^{B\times S\times d}\rightarrow\mathbb{R}^{BS\times d}$ and then compute the following metric which is a module level analogue of Wanda:
$\begin{split}\left(\rho\in\mathbb{R}^{d}\right)\propto\hat{\mathbf{a}}&=\frac{1}{o}\sum_{o}\mathbf{a}^{o}\\ \mathbf{a}^{o}&=\bigg{|}W[:,o]\bigg{|}\odot\mathrm{RootMeanSquare}\bigg{(}\mathbf{a},\mathrm{axis=}0\bigg{)}\end{split}$ (6)
Self-Attention Module: Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the self-attention module. For any data-sample sequence $b$, the output of the self attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output. We can flatten $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}\rightarrow\mathbb{R}^{BSd_{h}\times h}$ and then use the same formula as Equation 5 above to calculate $\rho\in\mathbb{R}^{h}$.
Appendix G How many perturbative samples are reasonable?
For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
Table 14: Varying the number of sub-models generated. Wikitext-2 Perplexity. LLaMA-2 7B pruned to 50% sparsity
Num Samples w/o Post-Pruning Adaptation w Post-Pruning Adaptation
1000 $22.09$ $9.25$
200 $61.63$ $9.15$
50 NaN $9.24$
Using $\mathrm{ns}_{\mathrm{sub-models}}=50$ results in an model with NaN perplexity on the Wikitext validation set. We posit that this is because of the LLaMA models are half precision, and removing the wrong modules can result in activations going outside of the FP16 dynamic range for unique data-points. Note that we are able to recover good performance of the model after fine-tuning though (we do not observe NaNs with the Wikitext-2 training data). This indicates that Bonsai actually recovers good modules even using as few samples as 50 sub-models.
Appendix H Mistral-7B Experiment Details
In addition to the primary experiments on the LLaMA and Phi-2 models, supplementary experiments were performed on the Mistral-7B (Jiang et al., 2023) model in comparison with Wanda results on the stated model. We apply Bonsai with the same hardware and configuration settings as used for the LLaMA and Phi-2 experiments. We target different pruning fractions (0.05, 0.1, and 0.2) across different numbers of samples and masks per iteration to evaluate the method’s performance under varying sparsity conditions.
The Mistral-7B model architecture differs from the LLaMA architecture in its use of group query attention and sliding window attention in lieu of the standard self-attention used in most transformer-based models like LLaMA (Jiang et al., 2023). We factor these differences into consideration in the implementation of Bonsai for Mistral. For the experiments that produced the results below, all Bonsai hyper-parameters are the same as Appendix A.2.
Table 15 presents the test perplexity results for Mistral-7B under different pruning methods. Considering the fully-structured sparsity nature of Bonsai, it achieves a test perplexity of 47.5 without post-pruning adaptation, with 1.66$\times$ inference speedup. After performing post-pruning adaptation on our pruned Mistral-7B, perplexity dropped drastically to 10.08. Note that the reported results of Wanda-pruned Mistral-7B are not fine-tuned afterward; if they were, their results would be marginally better than Bonsai’s results. However, as shown in Table 2, latency speedup would have dropped rapidly, while Bonsai stays the same at 1.66$\times$.
Table 15: Test perplexity of Mistral-7B model on Wikitext-2 across fully-structured Bonsai and semi-structured Wanda methods.
Sparsity Level Method Test PPL
Original, unpruned Mistral-7B N/A N/A 5.245
Wanda semi-structured 2-4 magnitude 13.81
Wanda 12.38
SparseGPT 10.46
Bonsai (w/o Adaptation) structured 50% magnitude 67.48
Wanda 47.50
Bonsai (w/ Adaptation) structured 50% Wanda 10.08
We further investigate the pruning habits of Bonsai by examining the pruned layers of Mistral, as shown in Figure 4. We notice a recurring theme: when an attention layer is significantly altered, it leads to compensation in the next layers within the sequence. This adaptive behavior, termed the ”Hydra effect” by McGrath et al. (2023), implies that the layers within a language model interact in a way that changes in one layer prompt adjustments in another. McGrath et al. (2023) specifically mentioned that when one attention layer was removed from a language model, the model was still able to self-repair and produce similar outputs; but it did so by relying more heavily on other layers.
Figure 4: Mistral’s pruned attention layers. The heavily pruned layers are usually preceded by or sandwiched between lightly-pruned layers, exhibiting the self-repairing ”Hydra effect” (McGrath et al., 2023).